OpenStudy (anonymous):

Solve the following equation. SHOW WORK! 5/8f=9

OpenStudy (anonymous):

are you solving for f?

OpenStudy (anonymous):

then start by multiplying 5 to both sides then it would be 8f=45 then divide 8 to both sides so f=5.(:

OpenStudy (anonymous):

not 5 hold on

OpenStudy (anonymous):

f=5.625(:

OpenStudy (lgbasallote):

divide both sides by 5/8 f = 9/5/8 get the reciprocal of 8 then multiply to 9/5 f = 9/5 times 1/8 f = 9/40